FFIR ISSUES TO BE STUDIED

FFIR group December 20, 1999

Contents

1 Parameter 4

2 Beam Delivery (switch/matching) System 4 2.1 Two IPs for γγ, γe^{−} collisions . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2 Two IPs for a concurrent e^{+}e^{−} collision . . . . . . . . . . . . . . . . . . . . 5

3 Collimation 5 3.1 collimation depth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 material . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 wakefield . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.4 muon background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Final Focus 7 4.1 Common optics for 0.3 - 1.5 TeV . . . . . . . . . . . . . . . . . . . . . . . 7 4.2 Tuning/Feedback method . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

5 IR 8 5.1 Horizontal Crossing Angle: Crab cavity? . . . . . . . . . . . . . . . . . . . 8 5.2 Final focus Quadrupole Magnet . . . . . . . . . . . . . . . . . . . . . . . . 8

5.2.1 configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 5.2.2 support system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

5.3 Superconducting compensation magnet . . . . . . . . . . . . . . . . . . . . 9 5.4 Detector Solenoid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5.5 Revised optics for larger clearance of SR . . . . . . . . . . . . . . . . . . . 9 5.6 Diagnostics? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5.7 Vacuum chamber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6 Detector and Background 9 6.1 Energy Range, 0.25 - 1 - 1.5 TeV? . . . . . . . . . . . . . . . . . . . . . . . 9 6.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6.2.1 synchrotron radiation . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6.2.2 muon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 6.2.3 e^{+}e^{−} pairs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 6.2.4 neutrons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 6.2.5 minijets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

6.3 Active mask and veto calorimeter inside the mask . . . . . . . . . . . . . . 10 6.4 Compact Detector as proposed for NLC . . . . . . . . . . . . . . . . . . . . 11

7 Luminosity Measurement 11 7.1 as a function of center-of-mass energy . . . . . . . . . . . . . . . . . . . . 11 7.2 beam size measurement by e^{±} pairs . . . . . . . . . . . . . . . . . . . . . . 11 8 Beam Extraction and Dump 11

8.1 recycle? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 8.2 measurement of beam polarization . . . . . . . . . . . . . . . . . . . . . . . 12 8.3 measurement of beam energy . . . . . . . . . . . . . . . . . . . . . . . . . 12

9 Instrumentation 12 9.1 magnet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 9.2 BPM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 9.3 BSM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 9.4 mover . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 9.5 control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 9.6 machine protection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

1 Parameter

considered.

Figure 1: JLC parameters in JLC Design Study, April 1997.

2 Beam Delivery (switch/matching) System

There are two interaction regions for two independent experiments. For minimum interference, the transverse and longitudinal separations are 20m and 200m, respectively, so that the optics is slightly different at left and right beam lines.

Figure 2: Schematic layout of the beam delivery system.

2.1 Two IPs for γγ, γe^{−} collisions

T.Takahashi(Hiroshima univ.)

2.2 Two IPs for a concurrent e^{+}e^{−} collision

Pulse by pulse switching?

3 Collimation

Non-linear optics is employed with sextupole magnets for 1.2km/0.75TeV(beam energy) collimator in order to selectively blow up the tail part of the beam. The collimated beam profile is 6σ_{x} × 40σ_{y}. For NLC, linear optics is employed for 2.4km/0.5TeV(beam energy) collimator. The collimated beam profile is 7σ_{x} × 35σ_{y}. Shorter collimation section is desired generally. New idea is welcome, and we may expect exotic ideas such as laser, liquid metal-collimation(?).

3.1 collimation depth

SLAC/NLC study shows 10^{−4} attenuation with a 2 X_{o} . Is it enough?

3.2 material

When the beam hits on the collimator, the temperature grows up on the surface. The energy deposit on the surface (J/g) can be expressed by,

W_{e} = N_{e} · \frac{dE}{dX} (ev · \frac{cm^{2}}{g}) · \frac{1.}{a} · 1.6 × 10^{−19}(\frac{J}{eV} )  (1)

where N_{e} particles hit on the surface area of a(cm^{2}). The thermal energy due to a temperature rise (δT) is given by,

W_{T} = 3 · \frac{6.02 × 10^{23}}{A} · 1.38 × 10^{−23}(\frac{J}{K}) · δT,  (2)

where, A is an atomic mass number. The temperature rise due to the energy deposit can be obtained by W_{T} = W_{e}. Apparently, a material with smaller atomic number is better for the collimator if its melting temperature is high enough. The relevant properties of several atoms are listed in table 1. Carbon seems to be the best on this point. Is it true?

Table 1: Candidate materials for collimator

material A  ρ  X_{0} T_{melt} T_{boil} T_{stress}
g/cm^{3} cm ^{∘}C ^{∘}C  ^{∘}C
C 12 2.25 18.8 3700 4800 < 2500
Ti  47.9 4.51 3.56 1660 3277 < 1000
W 183.8 19.24 0.35 3410 5900 ?
Cu  63.5 8.93 1.43 1080 3000 < 180
Fe  55.8 7.87 1.76 1535 2750  ?

3.3 wakefield

K.Yokoya derived a theoretical expression of wakefield induced by the collimators, which shall be compared to a measurement. The emittance growth due to the wakefield is expressed by,

Δε_{y} = (\frac{Δy^{2}}{β_{y}}) · (1.0 + \frac{Δ^{2}}{2.0}), Δ = 1.77Nr_{e} \frac{σ_{y}^{2}}{ε_{y}} [\sqrt{\frac{λ}{σ_{z}}} (\frac{1.0}{θa^{2}} + \frac{L}{a^{3}} ) + 0.69\frac{θ}{aσ_{z}} ] (3)

where λ ≡ 1.0/μ_{o}cρ = 1/120πρ is the resistive depth of material of the collimator, and ρ is the conductivity, and N, ε_{y}, σ_{y}, σ_{z} are the intensity/bunch, (vertical) invariant emittance, vertical beam size, bunch length of the beam, respectively. r_{e} is the electron classical radius. The geometrical parameters of the collimator are displayed in Fig.3. There is the optimal angle (θ_{opt}) minimizing the Δ, that is, (2λσ_{z})^{1/4}/a^{1/2}. With this angle and β_{y} instead of ε_{y} and σ_{y}, the Δ is,

Δ = 1.77Nr_{e} \frac{β_{y}}{γ} [\sqrt{\frac{λ}{σ_{z}}} \frac{L}{a^{3}} + \sqrt{2}\frac{θ_{opt}}{aσ_{z}} ]  (4)

where γ is the Lorentz factor, E_{beam}/m_{e}.

The emittance growth should be less than 1% so that the total growth due to 6 collimators is less than 6%. Assuming that the surface of the collimator is plated by copper, the aperture (a) of the collimator must be 250 (204)μm corresponding to 70.4 (56.9) σ_{y} with the length of 20 (10)cm, N = 10^{10}/bunch σ_{z}=80μm and β_{y}=400m at E_{beam} =750GeV, where the optimal angle (θ_{opt}) is 1.05^{o} (1.17^{o}) .

Figure 3: Geometry of a collimator.

3.4 muon background

(Y.Namito) It consists of 6 cylinders (iron or lead) with 0.6ϕ × 120m in each. For NLC, it consists of 4 spoilers as a tunnel filler of 3 × 3 × 9m^{3}.

Muon transformation is significantly sensitive to a detailed geometry of tunnel. Radioactivity must be studied inside/outside the tunnel. Optimization of the collimation method is necessary by using the two methods.

4 Final Focus

4.1 Common optics for 0.3 - 1.5 TeV

(K.Oide)[2] The basic idea of the optical system, consisting of two CCS(chromaticitycorrection section with two identical sextupole magnets connected by a pseudo-I transformer), the final ”telescope” and matching sections, is equivalent to that of the optics of FFTB. The total length of final focus system is 1600 m per beam for a common beam line at 0.3 < \sqrt{s} < 1.5TeV except for the final doublets which must be changed twice at least. The momentum acceptance is δp/p = ±1.0%, and the total beam intensity in a bunch-train must be very stable within ±0.5%. If the system is optimized at \sqrt{s}=0.5 (1.5)TeV, the length is 500 (1100) m. The above optimization should be discussed.

4.2 Tuning/Feedback method

(N.Yamamoto, S.Matsumoto)[1] Ground motion can be corrected at the frequency of less than 10Hz with BPMs by using beam-beam deflection, that is slow feedback of SLC type. The ground motion, whose amplitude is at the order of nm at > 10Hz, causes 5% loss of luminosity. BPMs with 10nm resolution is necessary for a beam based alignment (orbit correction).

For NLC, in addition, a fast feedback system is considered within 2.8nsec at <200 Hz by BPMs with pilot beam and also by beam-beam deflection. Feedback systems of SLC and B-factories can be again the prototype. The feedback simulation is necessary with 10-100nm resolution BPMs.

5 IR

5.1 Horizontal Crossing Angle: Crab cavity?

The design principle is the smallest crossing angle as possible because of the highest luminosity without a crab cavity. The present angle of ±4mrad is limited by synchrotron radiation background emitted from collimated beam. If we use a crab cavity, the luminosity increases by about 40%. For NLC, SLAC assumes the perfect performance for a crab cavity, then the crossing angle is ±10mrad for easier extraction of disrupted beams, whose magnitude is limited by a 4 Tesla (detector-)solenoidal field. Phase stability of the crab cavity must be less than 0.2^{∘}, thus we need a prototype cavity. Since KEK-B factory uses a crab cavity, it may be the prototype(?).

5.2 Final focus Quadrupole Magnet

5.2.1 configuration

The present design is a warm magnet (FeCo alloy) of 2.2 m long with the inner radius of 6.86mm and the pole tip field of 1.3 Tesla. It is located at ℓ^{∗}=2m from IP. For NLC, it consists of two permanent magnets of 1m long each, whose inner (outer) radii are 7(20) and 8(25) mm, respectively, and a superconducting quadrupole magnet(Q1SC) of 0.5 m long. PEP-II also uses permanent magnets. It is also located at ℓ^{∗}=2m from IP. For the case of the warm magnet, a water cooling must not cause any vibration. Beam-based alignment must be very difficult with the permanent magnets.

For the third option, a superconducting quadrupole magnet can be used with the smallest crossing angle and a clever extraction method of beams. This option is very attractive because no compensating magnet is necessary. So, it should be seriously considered. The parameters of the superconducting quad were given in the JLC-1[1], which are the aperture radius of 6.0cm, the inner radius of coil of 8.0cm, the outer radius of cryostat of 30cm, the field gradient of 117 Tesla/m and the material of NdTi at 1.8 K (superfluid He). The current stability must be less than 4 × 10^{−6}.

The longer distance(ℓ^{∗}) of the quad from IP makes a smaller dead cone and less background (back-scattered photons), and it must be benefit if it is set outside the compact detector. If ℓ^{∗}= 1 (3)m, the final focus system becomes shorter (longer) by 25 (20) % (NLC study).

5.2.2 support system

A simple simulation shows that a support tube is enough to stabilize relative positions of the final focus magnets at both sides of electron and positron beams with an input ground motion measured at TRISTAN-Fuji experimental hall[4]. For NLC, additional ”optical anchor” is considered, which is based on laser interferometer grounded on bed rock. SLD vibration was measured to be 30-50nm at f>1Hz. The laser interferometer can measure 1nm vertical oscillation of quadrupole magnets.

5.3 Superconducting compensation magnet

There must be superconducting solenoid magnets around the final quads in order to compensate the detector solenoidal field for the cases of warm and permanent magnets. Since the compensating magnet has a cryostat etc. , the permanent magnet option may loose one of the big advantages. A thinner cryostat is desired for the smaller dead cone.

5.4 Detector Solenoid

The present design is 2 Tesla. For NLC, 4 Tesla magnet was considered, and even higher magnetic field is desired for the compactness. However, the magnetic field must be optimized with a performance of calorimeter. The high field also limits the crossing angle.

5.5 Revised optics for larger clearance of SR

(K.Oide, T.Tauchi)

5.6 Diagnostics?

Shintake monitor at IP, better about 50cm around IP.

5.7 Vacuum chamber

T.Omori

6 Detector and Background

6.1 Energy Range, 0.25 - 1 - 1.5 TeV?

Snowmass96 etc.: 1.5TeV is necessary from SUSY-complementary to LHC. General questions are ; (1) does the proposed detector work for 0.25 < \sqrt{s} < 1.5TeV?, (2) does background constraint a detector type?, (3) detector upgrade plan for \sqrt{s} >1TeV ?

6.2 Background

6.2.1 synchrotron radiation

There is expected to be no serious background of synchrotron radiations because the aperture of the final quad is optimized with the collimated beam and the synchrotron radiations from upstream of last bend can be shielded by masks. However, SLD experiment observed a large fluctuation of background in the central drift chamber(CDC). What is expected to be a stability of beam with reasonable luminosity ? NLC considers synchrotron radiations from the 7σ_{x} × 35σ_{y} beam. The major source is the inner surface(4.5mm radius) of QA closest to the IP, length 1.8m, between 2.0 and 3.8m from the IP by EGS4 and GEANT.

6.2.2 muon

Muon Background; a detector is the IP within a radius of 4.5m from the bean line. Only ”tunnel-filler” spoilers (750t) meet the design goal. A compact detector is better because of smaller acceptance of muons.

6.2.3 e^{+}e^{−} pairs

The e^{+}e^{−} pairs produce 3.6 hits/mm^{2}/train at r=2.5cm (VTX). In the CDC, backscattered photons are converted to create 100 hits/train in total at \sqrt{s}=500GeV. The maximum allowable background level, traditionally: 1 hit/mm^{2}/train-crossing for a pixelbased vertex detector, and 10,000 photons/train-crossing for a drift chamber which corresponds to 1% occupancy with 1% conversion efficiency.

EGS4 results for 1TeV NLC shows 5-10 hits/mm^{2}/train at r=1cm and less than 1 hit/mm^{2} at r=2cm, however 10 times more hits by backscattered particles to be reduced by a low-Z mask. 30K photons/train penetrate the CDC at r=30cm even at 4T!, which may not be tolerable for a drift chamber. Most of photons come from the interaction of the pairs with the beam pipe and from backscattered from the surfaces which are closes in z to the IP, namely the luminosity monitor at z=190cm, the front face of the quadrupole doublet at 2m, and the septum of the rf shield, currently at 189cm. Synchrotron radiation backgrounds: at r=30cm, z=±100cm is 167K for 2T and 110K for 4T per train-crossing!

Anyway, we need a performance study on CDC and VTX under a condition of backgrounds by ABEL/JIM simulation.

6.2.4 neutrons

Neutrons are also produced at a rate of 10^{6}n/train by the pairs which interact with the pole tips of the final quads, where 0.13 neutron is produced per 1 GeV energy deposit. For NLC, the VTX has 3 × 10^{−3} n/cm^{2}/train corresponding to 10^{7} n/cm^{2}/year. Since a tolerable neutron yield is 3 × 10^{9} n/cm^{2}, they cause no serious problem.

However, neutron yields from the beam dump must be problem because of the huge energy deposit there. Without an efficient neutron absorber, the detector is seriously damaged. The detailed design is needed for the extraction line through the beam dump, so-called dump line.

6.2.5 minijets

6.3 Active mask and veto calorimeter inside the mask

It will be very important for SUSY-physics like as a single photon experiment. A typical minimum veto-angle is 50mrad. First performance study has been recently executed by using JIM simulation. The result can be read at URL of ¡A HREF=”http://acfahep.kek.jp/subg/ir/talks/tauchi.dec.1999.lmam.html”¿ here¡/A¿ .

6.4 Compact Detector as proposed for NLC

”Straw-man detector”: a compact five-layer tracking system using double-sided 5μmresolution silicon strips in B=4T. Beam pipes: 750μm thick beryllium, r=1cm near the IP and 500μm thick stainless steel for the rest. RF-shields: 200μm thick copper. Vertex detector: 4 layers of silicon+beryllium at the radii of 1.2, 2.7, 3.8 and 4.8cm with | cos θ| <0.9.

7 Luminosity Measurement

7.1 as a function of center-of-mass energy

A general question is ” How to measure the luminosity distribution within a beam energy spread of about 1%? This measurement is very important for toponium physics. There is a scheme to measure the luminosity distribution by precise measurement of the acollinearity angles of Bhabha scattering.

7.2 beam size measurement by e^{±} pairs

The scattering angles of the pairs can be measured by two sets of double discs of pixel device at ±1 m from IP. The beam aspect ratio can be obtained from the measurement of the azimuthal angles. At SLAC, there is a doubt for this possibility because of the huge background of synchrotron radiations etc. .

8 Beam Extraction and Dump

For NLC, there is a chicane in extraction beam line to separate the electron beam and photons (synchrotron, beamstrahlung, bremmstrahlung radiations) with a common dump. As mentioned before, the dump is a huge neutron sources for severe background.

Recently (Nov.1999) K. Yokoya summarized requirements for a dump line as follows;

(1) FF optics and available space: Final focus optics is based on the JLC design study and the horizontal crossing angle is ±4m rad. A beam line for ±10m from IP is occupied by QC1 and QC2. So, any optical element for a dump line must be located at the upstream of 10m from IP.

(2) Beam parameters: Beam energy is assumed to be 500GeV for the most severe case. However, a design for 250GeV beam energy must be done separately because of its larger emmitance. For the beam parameters, JLC-C shall be chosen for the largest emmitance, which are listed in Table 2.

After collisions, beam parameters were estimated by CAIN, and they are listed in Table 3.

(3) Beam diagnostic: At the dump line, there should be three measurements of (a) energy spectrum ( within a peak ), (b) polarization and (c) positrons in pair creation.

(4) Energy spectrum: For the measurement of beam energy, a bending magnet must be needed. The magnet should be vertical bend to prevent the positrons(c) from bending

Beam energy  500GeV
norm.emit. x/y 500/14 × 10^{−8}m
β_{x}/β_{y}  13mm / 0.2mm
σ_{z}  145μm

Table 2: Beam parameters at the final focus system.

norm.emit. x/y  1600/20 × 10^{−8}m
β_{x}/β_{y}  4mm / 0.15mm
α_{x}/α_{y}  1.5/0.0

( nonzero because of beam-beam lens. )

Table 3: Beam parameters after collisions.

(horizontally) into the final focus line. NLC uses a horizontal bend because of its large crossing angle.

(5) Focussing optics: Extracted (disrupted) beams must be focused in order to be collided with lasers (polarized laser and laser wire etc.) for (a) and (b). The Rayleigh length(Z) of the laser is chosen to be equal to the bunch length of electron beam, σ_{z}=150μm. Since the minimal spot size of the λ=1μm laser is 3.5μm, the beam spot sizes, σ_{x}, σ_{y}, must be the same, 3.5μm. Taking accout of about twice emmitance at 250GeV beam, β_{x} and β_{y} should be 0.4m and 30m, respectively.

For the measurement of energy spectrum (a) with 0.1% resolution, a vertical dispersion must be reqiured to be η_{y}=3.5mm, where the corresponding spatial spread is 3.6μm ( = 0.1%×η_{y} ) as the same as the laser spot size. An orbit slope should be zero there for the polarization measurement,η′_{y} = 0.

The above parameters must be optimized.

8.1 recycle?

An idea of recycle-use of beams is very interesting from the view point of dump-less and positron source.

8.2 measurement of beam polarization

8.3 measurement of beam energy

9 Instrumentation

N.Toge

9.1 magnet 9.2 BPM

H.Hayano

9.3 BSM

H.Hayano

9.4 mover 9.5 control 9.6 machine protection

References

[1] JLC Group: JLC-1 (Green Book), KEK Report 92-16, December 1992, see also http://www-jlc.kek.jp/JLC.proposal-j.html.

[2] JLC Design Study Group: JLC Design Study, KEK Report 97-1, April 1997, see also ftp://lcdev.kek.jp/pub/DesignStudy/.

[3] T.Tauchi, K.Yokoya and P.Chen, Part. Accel. 41, 29 (1993).

[4] S. Kanda, proceedings of workshop on physics and experiments with linear e^{+}e^{−} colliders, Waikoloa, Hawaii, 26-30 April 1993, p.758.